• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name ZAFAR AHMED

    Log entry time 12:08:27 on July 22, 2010

    Entry number 331652

    Followups:

    keyword=Change of pdestal for CR sum in PAW analysis for Moller

    I change the value of pedestal from 210.8 to 235.8 for histogram number
    ADC 9 of lg_spectra.kumac 
    
    In Figure run_14581_old_ped.gif I plotted lg_spectra with pedestal value
    210.8 and we can see that pedestal is shifted lit bit to the right. So I
    changed the pedestal value of histogram ADC 9 from 210.8 to 235.8 in the
    file pedes_14355.dat. I save the old file as pedes_14355_origional.dat
    
    In Fiugre run_14581_new_ped.gif I plotted lg_spectra with new value of
    pedestal and we can see that now pedestal is at zero channel.
    
    I will change it back to previous value of pedestal as soon I am finished
    with my analysis.
    


    A copy of this log entry has been emailed to: brads , rom , gen , romanip , glamazdi , zahmed



    Figure 1



    Figure 2